How do I solve the following systems of equations by using the addition method?
2x-y=3
x-2y=9

Answer by robertb(5830)

Multiply the top equation by -2: -4x+2y = -6.
Then add to the bottom equation: -4x+2y + x-2y = -6+9, or -3x = 3, or x = -1.
This makes the bottom equation -1-2y = 9, or -2y = 10, or y = -5.